Using the Pythagorean theorem:
Two jets pass directly over each other at 10am. One jet is traveling south at 480 km/h the other is traveling east at 900 km/h. How far are the two jets apart 2 hours later?

Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

Using the Pythagorean theorem:
Two jets pass directly over each other at 10am. One jet is traveling south at 480 km/h the other is traveling east at 900 km/h. How far are the two jets apart 2 hours later
---------------
From the crossing point, one jet is 480 km and the other is 900 km
The hypotenuse is sqrt(480^2 + 900^2)
=sqrt(230400 + 810000)
= sqrt(1040400)
= 1020 km

Answer by checkley77(12844)

a^2+b^2=c^2
a=480*2=960
b=900^2=1800
960^2+1,800^2=c^2
921,600+3,240,000=c^2
c^2=4,161,600
c=2,040 km apart after 2 hours.